• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reimer

    Log entry time 18:15:34 on March18,2006

    Entry number 167363

    This entry is a followup to: 167358

    Followups:

    keyword=Compton lock

    MCC noticed that "compton lock" was not on. I told tham that is should be on, they asked for the beam position. Unfortunately, in http://hallaweb.jlab.org/compton/OSP/Data_taking/index.php there are two values given for BPM IPM102B: either -0.92 +/- 0.03 and -0.88. MCC tuned to -0.92 for now. The actual value is 0.94 righ now, but the rates in the photon and electron counters have improved. it is now possible to see the difference between laser on and laser off running.


    FIGURE 1